### v1

# APCTP-Pre2025-021

APCTP-Pre2025-021
Weakly model-independent determination
of total expansion during inflation
Dayeong Choia, [^1] Equal contribution, Subin Jeona,1 and Jinn-Ouk Gonga,b, [^2] Corresponding author
aDepartment of Science Education, Ewha Womans University, Seoul 03760, Korea
bAsia Pacific Center for Theoretical Physics, Pohang 37673, Korea
We study systematically the total expansion experienced by a certain perturbation mode during single-field inflation, not resorting to explicit models of inflation or reheating. By assuming that during the reheating stage the equation of state $w_{\text{rh}}$ can be written as a function of $e$-folds, the unknown dynamics during reheating parametrized by $w_{\text{rh}}$ is confined within a time integral so that any dependence on the models of inflation and reheating is isolated from model-independent contributions. Especially, the dependence on the reheating dynamics via $w_{\text{rh}}$ and the reheating temperature $T_{\text{rh}}$ is dominating. We give two illustrative examples of $w_{\text{rh}}$ to discuss its impacts on the total expansion, which can be different as much as 10 even for the same reheating temperature, depending on the shape of $w_{\text{rh}}$.
1 Introduction
The phase of accelerated expansion of the universe at very early times, cosmic inflation, provides homogeneity and isotropy of the observed universe as can be seen from the cosmic microwave background (CMB) [1, 2, 3]. At the same time, during inflation tiny quantum vacuum fluctuations are generated and become the primordial perturbations. After inflation, these small inhomogeneities are responsible for structure formation by gravitational instability [4, 5, 6, 7, 8]. The properties of these primordial perturbations have been constrained over decades from the observations on the temperature anisotropies of the CMB and the large-scale distribution of galaxies, and are consistent with the inflationary predictions [9].
To derive inflation, viz. to make the expansion of the universe accelerating, we need a very unusual matter content that has a strong negative pressure. One such a matter content is the cosmological constant, which however does not lead to a graceful exit from inflation. Instead, we usually introduce a scalar field whose potential energy is dominant over the kinetic one so that it behaves effectively as a cosmological constant, leading to an accelerated, quasi-exponential expansion of the universe. This scalar field that dominates the energy density of the universe during inflation is called the inflaton [10, 11, 12].
Meanwhile, the standard paradigm for physical cosmology, the hot big bang model, describes successfully the evolution of the universe starting from big bang nucleosynthesis (see e.g. [13, 14]) once the initial conditions consistent with current observations – homogeneous and isotropic universe with ${\cal O}(10^{-5})$ fluctuations – are provided by inflation. This raises a very serious question. During inflation, the energy density of the universe is dominated by the inflaton. When the standard hot big bang evolution commences after inflation, relativistic matter contents such as photons occupy most of the energy density of the universe, along with tiny fraction of pressureless components like baryons and dark matter. This means the energy of the inflaton sector must be transferred to the usual matter contents with non-negative pressure, filling the universe with energetic particles. This process of energy transfer is called reheating (for reviews, see e.g. [15, 16, 17]).
Thus, without reheating, after inflation the universe never becomes hot and thermalized, not leading to the universe as observed. Reheating is an indispensable epoch that connects inflation and hot big bang. But as the universe, after reheating, is in thermal equilibrium, any model-dependent features that would have characterized the details of the reheating mechanism are swept away. This makes it very difficult to study the epoch of reheating. All we know is that the inflaton should decay and/or annihilate, and the final products are the standard model particles and dark matter. A standard, old approach is to introduce a decay width of the inflaton so that reheating proceeds perturbatively [18, 19, 20]. However, since reheating epoch is so elusive, any non-perturbative process may well take place, such as preheating [21, 22], rescattering [23, 24] and turbulence [25, 26, 27].
Connecting the reheating mechanism directly to the preceding inflationary epoch demands all the detailed properties of the inflaton – how it is coupled to other species, what is the form of the potential after inflation around the minimum, and so on. However, as the identity of the inflaton itself is unknown, its properties are completely behind the veil. Thus an alternative approach is taken. Considering the simple inflation model with a quadratic potential, after inflation around the minimum of the potential the rapid oscillation of the inflaton makes it behave effectively as a pressuress matter, characterized by the equation of state $w\equiv p/\rho=0$. Thus, phenomenologically we can adopt a specific equation of state during the reheating stage that encrypts a certain reheating mechanism [28].
In this article, we study systematically the contributions to the total expansion during inflation of a certain perturbation mode experiences. These contributions can be categorized into those dependent on and independent from the models of inflation and reheating dynamics. Only assuming that the equation of state during reheating can be written as a function of $e$-folds, we show that all the model dependence can be isolated from the model-independent contributions. This article is organized as follows. In Section 2, we consider the total expansion of a mode and separate individual contributions. By assuming that the equation of state during reheating is a function of $e$-folds, model-dependent contributions are sequestered from the model-independent ones. In Section 3 we study the two model-dependent inputs, the endpoint of inflation and the equation of state during reheating. We present a more accurate estimate on the endpoint given an inflation model, along with a simple, more phenomenological parametrization. We give explicit examples for the equation of state during reheating, and show the total expansion of a mode as a function of model parameters explicitly. Finally we conclude in Section 4.
2 Total expansion a mode experiences
2.1 Contributions to total expansion
We begin with the ratio of comoving Hubble scale $a_{k}H_{k}=k$, when a certain perturbation mode with comoving wavenumber $k$ exited the horizon during inflation, to the present one $a_{0}H_{0}$. Here, we follow the minimal assumption that right after inflation the reheating stage follows. After reheating is complete the evolution of the universe can be described by the standard cosmological model including radiation and matter dominated epochs. Then, we may write the ratio $k/(a_{0}H_{0})$ as [29]:
$\frac{k}{a_{0}H_{0}}=\frac{a_{k}H_{k}}{a_{0}H_{0}}=\frac{a_{k}}{a_{e}}\frac{a_{e}}{a_{\text{rh}}}\frac{a_{\text{rh}}}{a_{0}}\frac{H_{k}}{H_{0}}\,.$ (1)
Taking logarithm of both sides, and shuffling some terms gives
$N_{k}=-\log\bigg(\frac{k}{a_{0}H_{0}}\bigg)-N_{\text{rh}}+\log\bigg(\frac{a_{\text{rh}}}{a_{0}}\bigg)+\log\bigg(\frac{H_{k}}{H_{0}}\bigg)\,$, (2)
where $N_{k}=\log(a_{e}/a_{k})$ and $N_{\text{rh}}=\log(a_{\text{rh}}/a_{e})$ are the numbers of $e$-folds respectively from the moment during inflation the mode with comoving momentum $k$ crosses the horizon till the end of inflation, and from the end of inflation till radiation dominated epoch begins, viz. the end of reheating stage.
  1. 1.
    $N_{k}$: To explain the observed homogeneity and isotropy of CMB, we demand that this number be larger than a certain value in any successful inflation model. Assuming that inflation is driven by a single inflaton field $\phi$, $N_{k}$ can be estimated by trading time with the value of $\phi$ as
    $N_{k}=\int_{k}^{e}Hdt=\int_{\phi_{k}}^{\phi_{e}}\frac{H}{\dot{\phi}}d\phi\,$, (3)
    where we have assumed that $\phi=\phi(t)$. Thus, to determine $N_{k}$, we need two inputs. On one hand, we need to specify exactly the two boundary values of $\phi$: $\phi_{k}$ when the $k$-mode crosses the horizon, and $\phi_{e}$ when inflation ends. On the other hand, we should also be able to write both $H$ and $\dot{\phi}$ as functions of $\phi$. However, it is usually impossible to find the full analytic solution for $\phi(t)$ as we need to solve the highly non-trivial equation of motion for $\phi$. Instead, we usually resort to the slow-roll approximation to gain analytic control. Furthermore, solving for $\phi(t)$ to determine two boundaries and to find $H$ and $\dot{\phi}$ means that we specify the model of inflation.
  2. 2.
    $\log\big[k/(a_{0}H_{0})\big]$: Since $a_{0}$ is arbitrarily normalized, conventionally we set $a_{0}=1$. Then, the reference wavenumber $k$ corresponds to the present value. Regarding the CMB observations, $k={\cal O}(10^{-3}-10^{-2})$ Mpc-1. The present Hubble parameter $H_{0}$ can be determined by various observations on, for example, CMB [9]. Thus this term can be pinned down without any ambiguity up to observational errors in $H_{0}$.
  3. 3.
    $N_{\text{rh}}$: This number measures the expansion of the universe starting from the end of inflation until reheating ends so that the universe becomes dominated by radiation. As $N_{\text{rh}}$ depends on the detail of the reheating mechanism that remains mostly elusive, we usually resort to an effective and/or phenomenological approach. One possible way is to make use of the equation of state during reheating, $w_{\text{rh}}$, and to assume that it is a constant, e.g. $w_{\text{rh}}=(n-2)/(n+2)$ for $V(\phi)\propto\phi^{n}$ when averaged over oscillation cycles [30]. This leads to the following relation between the energy densities at the end of inflation $\rho_{e}$ and at the end of reheating stage $\rho_{\text{rh}}$:
    $N_{\text{rh}}=\frac{1}{3(1+w_{\text{rh}})}\log\bigg(\frac{\rho_{e}}{\rho_{\text{rh}}}\bigg)\,.$ (4)
  4. 4.
    $\log(a_{\text{rh}}/a_{0})$: Since radiation domination, we can follow the standard thermal history of the universe. Assuming entropy conservation after reheating, $a_{\text{rh}}/a_{0}$ can be written as (see e.g. [31])
    $\frac{a_{\text{rh}}}{a_{0}}=\bigg(\frac{11}{43}g_{*S}\bigg)^{-1/3}\frac{T_{0}}{T_{\text{rh}}}\,$, (5)
    where $g_{*S}$ is the effective number of light species for entropy at the moment of reheating. Thus, in determining $\log(a_{\text{rh}}/a_{0})$, the most unclear factor is the reheating temperature $T_{\text{rh}}$. This is closely related to $N_{\text{rh}}$, as once it is figured out we can calculate immediately, starting from $\rho_{e}$, the energy density at the end of reheating $\rho_{\text{rh}}$, which gives $T_{\text{rh}}$ by
    $\rho_{\text{rh}}=\frac{\pi^{2}}{30}g_{*}T_{\text{rh}}^{4}\,$, (6)
    where $g_{*}$ is the effective number of relativistic species at the moment of reheating, and is not necessarily the same as $g_{*S}$. Both $g_{*}$ and $g_{*S}$ depend on temperature, and assuming the validity of the standard model of particle physics up to ${\cal O}(100)$ GeV, $g_{*}=106.75$.
  5. 5.
[…]
2.2 Sequestering model dependence
As we have seen, estimating $N_{k}$ exactly requires that we specify the model of inflation and solve the equations numerically. Even if we are to make use of the right-hand side of (2), we still need to know the model of inflation via $\rho_{e}$ and $H_{k}$, and additionally, on top of inflation, the detail of the reheating process through $N_{\text{rh}}$ and $a_{\text{rh}}$. Thus one may be tempted to conclude that at least without specifying the inflation model it is difficult to estimate $N_{k}$. But we can improve the situation as follows. If the universe is dominated by a canonical scalar field, viz. the inflaton, the acceleration of the scale factor is determined by the following equation, with $\rho_{\phi}=\dot{\phi}^{2}/2+V$ and $p_{\phi}=\dot{\phi}^{2}/2-V$:
$\frac{\ddot{a}}{a}=-\frac{\rho_{\phi}+3p_{\phi}}{6m_{\rm Pl}^{2}}=\frac{\dot{\phi}^{2}-V}{3m_{\rm Pl}^{2}}\,.$ (7)
During slow-roll phase, $\dot{\phi}^{2}\ll V$ so that $\ddot{a}>0$, i.e. the expansion is accelerating and the universe undergoes inflation. Meanwhile, inflation ends at $\phi=\phi_{e}$ when the following condition is satisfied so that $\ddot{a}=0$:
$\dot{\phi}_{e}^{2}=V(\phi_{e})\,.$ (8)
Thus our first observation is that, from (8), at the end of inflation $w_{e}=-1/3$:
$w_{e}=\frac{p_{\phi}}{\rho_{\phi}}\bigg|_{e}=\frac{\dot{\phi}^{2}/2-V}{\dot{\phi}^{2}/2+V}\bigg|_{e}=-\frac{1}{3}\,.$ (9)
At the beginning of radiation domination, viz. at the end of reheating epoch, the equation of state is that of radiation: $w_{r}=1/3$. That is, at both ends of the reheating stage, the values of the equation of state are fixed, 1/3-1/3 at the beginning and $1/3$ at the end. Thus we expect that any realistic equation of state during reheating begins from 1/3-1/3, changes smoothly, and ends up at the value of 1/31/3.333Here, we assume that during the reheating state $w_{\text{rh}}$ does not become smaller than 1/3-1/3. If so, until $w_{\text{rh}}>-1/3$ again the universe undergoes yet another inflationary stage. Now we assume that $w_{\text{rh}}$ can be written as a function of $e$-folds, with $N=0$ at the beginning and $N=N_{\text{rh}}$ at the end of the reheating stage. Then $w_{\text{rh}}(N=0)=-1/3$ and $w_{\text{rh}}(N=N_{\text{rh}})=1/3$. From the continuity equation, we can solve for the energy density at the end of reheating, $\rho_{\text{rh}}$, with a general equation of state during reheating which is a function of the normalized $e$-folds $n\equiv N/N_{\text{rh}}$:
$\rho_{\text{rh}}=\rho_{e}\exp\left\{-3N_{\text{rh}}\int_{0}^{1}\big[1+w_{\text{rh}}(n)\big]dn\right\}\,.$ (10)
Thus, we can isolate the effects of the model-dependent equation of state during reheating within the time integration. On general ground, the integration of $w_{\text{rh}}(n)$ can be performed analytically and/or numerically:
$\int_{0}^{1}[1+w_{\text{rh}}(n)]dn=1+\int_{0}^{1}w_{\text{rh}}(n)dn\equiv 1+{\cal W}(\alpha)\,$, (11)
where $\alpha$ collectively denotes a set of parameters on which the evolution of $w_{\text{rh}}(n)$ depends. We expect that ${\cal W}(\alpha)$ is likely to be a number of ${\cal O}(1)$. We will consider explicit examples of $w_{\text{rh}}(n)$ later.
Now, using (5), we can write (2) as
$N_{k}=-\log\bigg(\frac{k}{a_{0}H_{0}}\bigg)-N_{\text{rh}}-\frac{1}{3}\log\bigg(\frac{11}{43}g_{*S}\bigg)+\log\bigg(\frac{T_{0}}{H_{0}}\bigg)+\log\bigg(\frac{H_{k}}{T_{\text{rh}}}\bigg)\,.$ (12)
Further, using (6), (10) and (11), the number of $e$-folds during reheating can be written as
$N_{\text{rh}}$ $=\frac{1}{3(1+{\cal W})}\log\bigg(\frac{\rho_{e}}{\rho_{\text{rh}}}\bigg)$
$=\frac{1}{3(1+{\cal W})}\bigg[\log\bigg(\frac{\rho_{e}}{m_{\rm Pl}^{4}}\bigg)-\log\bigg(\frac{\pi^{2}}{30}g_{*}\bigg)-4\log\bigg(\frac{T_{\text{rh}}}{m_{\rm Pl}}\bigg)\bigg]\,.$ (13)
Note that $N_{\text{rh}}\geq 0$ gives the upper bound on $T_{\text{rh}}/m_{\rm Pl}$ by demanding that the terms in the square brackets of (2.2) be non-negative. [^4] As long as we assume $w_{\text{rh}}>-1/3$, the prefactor $1+{\cal W}$ is always positive. For the extreme case of $w_{\text{rh}}=-1/3$ until the end of reheating, ${\cal W}=-1/3$. $T_{\text{rh}}$ is saturated when $N_{\text{rh}}=0$, viz. instantaneous reheating. Meanwhile, from the slow-roll approximation which is likely to be very effective for the scales corresponding to the CMB observations, we can write the power spectrum of the scalar perturbation as
${\cal P}_{\cal R}=\bigg(\frac{H}{2\pi}\bigg)^{2}\bigg(\frac{H}{\dot{\phi}}\bigg)^{2}=\frac{2}{\pi^{2}r}\frac{H^{2}}{m_{\rm Pl}^{2}}\,$, (14)
[…]
$\frac{H_{k}}{m_{\rm Pl}}=\sqrt{\frac{r{\cal P}_{\cal R}}{2}}\pi\,.$ (15)
[…]
Combining (12), (2.2) and (15), we find $N_{k}$ as
$N_{k}$ $=-\log\bigg(\frac{k}{a_{0}H_{0}}\bigg)-\frac{1}{3}\log\bigg(\frac{11}{43}g_{*S}\bigg)+\log\left(\frac{T_{0}}{H_{0}}\right)+\log\left(\sqrt{\frac{r{\cal P}_{\cal R}}{2}}\pi\right)$
$\quad+\frac{1}{3(1+{\cal W})}\log\bigg(\frac{\pi^{2}}{30}g_{*}\bigg)-\frac{4}{3(1+{\cal W})}\log\Bigg(\frac{\rho_{e}^{1/4}}{m_{\rm Pl}}\Bigg)+\frac{1-3{\cal W}}{3(1+{\cal W})}\log\bigg(\frac{T_{\text{rh}}}{m_{\rm Pl}}\bigg)\,.$ (16)
Notice that the terms in the first line of (2.2) can be all completely fixed within our observational errors and theoretical uncertainties, and are nearly independent of the detailed dynamics during inflation and reheating. Meanwhile, the terms in the second line of (2.2) – especially the last two terms – contain all the model dependence of inflation and reheating via ${\cal W}(\alpha)$, $\rho_{e}$ and $T_{\text{rh}}$. Note that if ${\cal W}=1/3$, the last term vanishes and $T_{\text{rh}}$ does not affect $N_{k}$. As can be read from (11), this is the case if simply $w_{\text{rh}}=1/3$, i.e. during reheating the universe is dominated by radiation. But even if $w_{\text{rh}}\neq 1/3$, we may say the universe is effectively radiation-dominated in the sense that ${\cal W}=1/3$ so that $N_{k}$ is independent of $T_{\text{rh}}$.
3 Model-dependent inputs
3.1 Endpoint of inflation
The energy density at the end of inflation, $\rho_{e}$, is certainly model-dependent. Usually, it is identified as the value of the potential energy when the potential slow-roll parameter $\epsilon_{V}\equiv m_{\rm Pl}^{2}(V^{\prime}/V)^{2}/2$ with $V^{\prime}=dV/d\phi$ becomes 1. For a given model of inflation, however, we can determine the endpoint of inflation without resorting to the slow-roll approximation as follows. Assuming no interaction with any other matter contents during inflation, $\phi$ satisfies the following equation of motion:
$\ddot{\phi}+3H\dot{\phi}+V^{\prime}=0\,.$ (17)
Here, we can define the second slow-roll parameter $\eta$ in such a way that
$\eta\equiv\frac{\dot{\epsilon}}{H\epsilon}=-\frac{\ddot{H}}{H^{3}\epsilon}+2\epsilon\,$, (18)
so we can replace $\ddot{\phi}$ with $\dot{\phi}$ and slow-roll parameters:
$\ddot{\phi}=-\frac{m_{\rm Pl}^{2}}{\dot{\phi}}\ddot{H}=-\frac{m_{\rm Pl}^{2}}{\dot{\phi}}H^{3}\epsilon(2\epsilon-\eta)\,$, (19)
and (17) can be written as
$\frac{m_{\rm Pl}^{2}}{\dot{\phi}}H^{3}\epsilon(2\epsilon-\eta)+3H\dot{\phi}+V_{\phi}=0\,.$ (20)
Now, at $\phi_{e}$ where (8) is satisfied, we can write $\dot{\phi}_{e}$ in terms of $V(\phi_{e})\equiv V_{e}$ as
$\dot{\phi}_{e}=-\sqrt{V_{e}}\,$, (21)
where we have assumed that during inflation $\phi>0$ and it approaches to 0, hence $\dot{\phi}<0$. Then from (20) we can find that at $\phi_{e}$ the following relation is exactly satisfied:
$\frac{V_{e}}{V_{e}^{\prime}}=\frac{2\sqrt{2}}{4+\eta_{e}}m_{\rm Pl}\,$, (22)
where $\eta_{e}$ is the value of $\eta$ at $\phi_{e}$ and is typically a constant of ${\cal O}(1)$. To have a more concrete idea, let us consider a general power-law potential, $V(\phi)=m^{4-n}\phi^{n}$. This gives directly
$\frac{\phi_{e}}{m_{\rm Pl}}=\frac{2\sqrt{2}n}{4+\eta_{e}}\,.$ (23)
The usual endpoint based on the potential slow-roll parameter $\epsilon_{V}\big|_{e}=1$ is
$\frac{\phi_{e}}{m_{\rm Pl}}=\frac{n}{\sqrt{2}}\,.$ (24)
In Table 1 for the power-law potential we compare the numerical value of $\phi_{e}/m_{\rm Pl}$ with (23) for different values of $\eta_{e}$, along with the usual slow-roll approximation result (24). Even if including interactions with other species, e.g. perturbative decay of $\phi$ as
$\ddot{\phi}+3H\dot{\phi}+V_{\phi}=-\Gamma\dot{\phi}\,$, (25)
with $\Gamma$ being a constant decay width, the value of $\phi_{e}$ is not changed appreciably as can be seen in Table 1.
$n$ 2 3 4
Numerical values 1.00938 1.66602 2.33939
(23) ($\eta_{e}=0.5$) 1.25708 1.88562 2.51416
(23) ($\eta_{e}=1.0$) 1.13137 1.69706 2.26274
(23) ($\eta_{e}=1.5$) 1.02852 1.54278 2.05704
(24) 1.41421 2.12132 2.82843
$\Gamma=0.1m$ 0.97188 1.66593 2.33939
$\Gamma=0.01m$ 1.00556 1.66601 2.33939
$\Gamma=0.001m$ 1.00899 1.66602 2.33939
Table 1: Comparison of $\phi_{e}/m_{\rm Pl}$ between numerical results, (23) with different values of $\eta_{e}$, and the standard slow-roll approximation (24) based on the power-law potential $V(\phi)=m^{4-n}\phi^{n}$. Typically $\eta_{e}=1$ gives reasonably good approximation for $\phi_{e}$. We also present $\phi_{e}/m_{\rm Pl}$ obtained numerically with different values of the decay width $\Gamma$ as given by (25).
[…]
$\log\left(\frac{\rho_{e}^{1/4}}{m_{\rm Pl}}\right)=\log\left(\beta^{-1/4}\frac{\rho_{k}^{1/4}}{m_{\rm Pl}}\right)=-\frac{1}{4}\log\bigg(\frac{\beta}{3}\bigg)+\frac{1}{2}\log\left(\sqrt{\frac{r{\cal P}_{\cal R}}{2}}\pi\right)\,.$ (26)
Note that we do not define $\beta$ as the ratio of $\rho_{\text{rh}}$ to $\rho_{e}$ as in [36], because then $\beta$, $N_{\text{rh}}$ and $T_{\text{rh}}$ are interchangeable. The constant $\beta$ as defined here characterizes a large class of inflation models. For example, for a large-field power-law potential model typically we have $\beta\gtrsim{\cal O}(100)$, while for hybrid inflation or hilltop inflation $\beta={\cal O}(1)$. However, as the dependence on $\beta$ is logarithmic, the contributions to $N_{k}$ of different models of inflation are not very different: The first term of (26), which in this way solely contains the dependence on the inflation model, gives 0.300993-0.300993, 0.876639-0.876639 and 1.45229-1.45229 for $\beta=10$, 100 and 1000 respectively.
3.2 Equation of state during reheating
As mentioned, we do not have any standard model of reheating process. So usually we have to make a set of reasonable assumptions and take phenomenological approaches to reheating. Here, we only assume that $w_{\text{rh}}$ can be written as a function of the normalized $e$-folds during reheating $n\equiv N/N_{\text{rh}}$, which encodes the detailed process during reheating as a smooth function starting from 1/3-1/3 and ending at $1/3$. As one example for such an equation of state during reheating, we model $w_{\text{rh}}(n)$ as the following:
$w_{\text{rh}}(n)=\frac{1-\alpha^{2}\big(n^{-1}-1\big)^{2}}{3\Big[1+\alpha^{2}\big(n^{-1}-1\big)^{2}\Big]}\,$, (27)
where $\alpha>0$ is a constant. In the left panel of Figure 1, we present the evolution of $w_{\text{rh}}$ for several different values of $\alpha$. Further, we can perform the integral (11) analytically to find ${\cal W}(\alpha)$ as
${\cal W}(\alpha)=\frac{1}{3(1+\alpha^{2})^{2}}\Big[(1-\alpha^{4})-\alpha\pi(1-\alpha^{2})-4\alpha^{2}\log\alpha\Big]\,.$ (28)
For $\alpha=0$, the value of $\cal W$} is 1/3, monotonically decreases and approaches 1/3-1/3 for $\alpha\to\infty$. Note that for $\alpha=1$, $w_{\text{rh}}(n)$ is an odd function around $n=1/2$ thus ${\cal W}=0$.
As another example, consider the following $w_{\text{rh}}(n)$:
$w_{\text{rh}}(n)=\frac{1}{6}\bigg\{\tanh\bigg[\alpha\bigg(\log\bigg(\frac{n}{1-n}\bigg)-\alpha\bigg)\bigg]+\tanh\bigg[\alpha\bigg(\log\bigg(\frac{n}{1-n}\bigg)+\alpha\bigg)\bigg]\bigg\}\,$, (29)
where $\alpha>0$. We show (29) as a function of $n$ in the middle panel of Figure 1 for different values of $\alpha$. In this case, since $w_{\text{rh}}(n)$ is symmetric around $n=1/2$, simply ${\cal W}=0$ irrespective of the value of $\alpha$. Then during the reheating epoch the energy density scales as that of pressureless matter, although the behaviour of $w_{\text{rh}}$ could be substantially different from $w_{m}=0$.
For these two examples, $w_{\text{rh}}(n)$ remains between 1/3-1/3 and $1/3$. But this needs not be the case if the universe after inflation experiences an exotic period. For example, during kinetic domination the equation of state becomes 1. As such an example with $w_{\text{rh}}\geq 1/3$, we consider the following:
$w_{\text{rh}}(n)=\frac{2n-1}{3\big(2n^{2}-2n+1\big)}+\frac{\alpha}{3}\exp\bigg[-\log^{2}\bigg(\frac{1-n}{n}\bigg)\bigg]\,.$ (30)
As shown in the right panel of Figure 1, (30) exhibits at $n=1/2$ a bump the height of which is set by $\alpha$ and can exceeds $1/3$. We found it is difficult to perform the integral of (30) analytically, but numerically it is easily done and $\cal W$} is linearly proportional to $\alpha$, with ${\cal W}(\alpha=0)=0$.
Figure 1: Plots of $w_{\text{rh}}$ as a function of $n\equiv N/N_{\text{rh}}$ given respectively by (left) (27), (middle) (29) and (right) (30). Here, (solid) $\alpha=0.1$, (dot-dashed) 0.5, (dense dot) 1, (sparse dot) 2 and (dashed) 5 respectively.
One may, by taking into account the oscillating inflaton field around the minimum of the potential, include sinusoidal modulations that diminish at both ends into $w_{\text{rh}}$ as the following:
[…]
Such a (slightly) oscillating behavior might be a more realistic approximation for $w_{\text{rh}}$. However, integrating such oscillations over the reheating period typically vanishes. This leads to little change in $\cal W$} even if we include sinusoidal modulations. Thus, we do not consider any diminishing oscillations in $w_{\text{rh}}$, not because such oscillations are unrealistic but because they do not lead to any appreciable difference.
Thus, as examples for a non-trivial equation of state during reheating epoch, we take (27) and (30) and see the corresponding number of $e$-folds a certain $k$-mode has gained during inflation. For the model-independent contributions, we take the following values from observations and theoretical estimates based on the standard model of particle physics:
[…]
1 $67.4\text{ km s}^{-1}\text{Mpc}^{-1}$ 106.75 2.725 K $2.0968\times 10^{-9}$
Table 2: Input values for the model-independent contributions in (2.2).
As a fiducial value of the tensor-to-scalar ratio, we take $r=10^{-3}$. We consider two representative wavenumbers, $k=0.05$ Mpc-1 and $k=0.002$ Mpc-1. The model-independent contributions in (2.2), viz. the first four terms, give 48.1115 for $k=0.05$ Mpc-1, and 51.3304 for $k=0.002$ Mpc-1 respectively. Then, for a given model of inflation, i.e. for a fixed value of $\beta$, (2.2) with (26) and $w_{\text{rh}}(n)$ becomes a function of $\alpha$ and $T_{\text{rh}}$.
Figure 2: Plots of $N_{k}$ for (upper panels) (27) and (lower panels) (30) as a function of (horizontal axis) $T_{\text{rh}}/m_{\rm Pl}$ and (vertical axis) $\alpha$. In both upper and lower panels, the reference scale is set to be (upper rows) $k=0.05$ Mpc-1 and (lower rows) $k=0.002$ Mpc-1. The value of $\beta$ is set to be (left panels) $\beta=1.5$, (middle panels) $\beta=10$ and (right panels) $\beta=100$. $T_{\text{rh}}/m_{\rm Pl}$ is presented up to the value that gives $N_{\text{rh}}=0$.
In Figure 2, we show (2.2) as a function of $T_{\text{rh}}/m_{\rm Pl}$ and $\alpha$ for given values of $\beta$, with the equation of state during inflation being (upper panel) (27) and (lower panel) (30). In the upper panels where we have considered (27), the larger $\alpha$ is, the smaller $N_{k}$ is. This is because, as can be read from the left panel of Figure 1, a larger value of $\alpha$ gives a slower progress of reheating, giving rise to a lower value of $N_{k}$. Note that as $\alpha\to 0$, it is no different from instantaneous reheating so there is no constraint on $N_{k}$ from the possible value of $T_{\text{rh}}/m_{\rm Pl}$. But generally, depending on the value of $\alpha$, even for the same reheating temperature $N_{k}$ could be different as large as ${\cal O}(10)$. In the lower panels where the equation of state during reheating is taken as (30). As the equation of state exceeds $1/3$ for a larger value of $\alpha$, we then have a larger values of $N_{k}$. This happens when ${\cal W}=1/3$ so that effectively $\cal W$} is identical to the case of instantaneous reheating. This occurs at $\alpha\approx 2.51171$ where, as can be seen, $N_{k}$ is independent of $T_{\text{rh}}/m_{\rm Pl}$. Across this line $\Delta{N}_{k}={\cal O}(10)$ for a wide range of $\alpha$.
4 Conclusions
In this article, we have studied the total expansion a certain mode experiences during inflation since the moment of horizon crossing until the end of inflation. In doing so, we have shown that the contributions which are dependent on the detailed dynamics of inflation and reheating are separated from those which are not. Furthermore, by assuming that the equation of state during the reheating stage $w_{\text{rh}}$ is a function of $e$-folds, its effects, i.e. the speculative dynamics during reheating, are confined within a time integral, giving rise to a simple function of model-dependent parameters with magnitude ${\cal O}(1)$.
We have illustrated the total expansion of certain $k$-modes by two choices of the equation of state during reheating, (27) and (30). Depending on the parameter $\alpha$ that controls the shape of $w_{\text{rh}}$ and the reheating temperature $T_{\text{rh}}$, the number of $e$-folds can be specified experienced by the $k$-mode. We could also confirm the two well-known observations that 1) slower reheating gives a smaller value of $N_{k}$ and, conversely, 2) a larger value of $N_{k}$ is obtained for $w_{\text{rh}}\geq 1/3$.
Therefore, by an appropriate modeling of the equation of state during reheating as a function of $e$-folds and by choosing the desired reheating temperature, the impacts of the unknown dynamics of reheating on the expansion of perturbation modes can be easily estimated. For any generic models of reheating it is possible to parametrize $w_{\text{rh}}$ as a function of $e$-folds even if we cannot derive the analytic form of $w_{\text{rh}}$ microscopically, e.g. by numerical fitting. Thus, we believe our approach can be universally adopted for generic studies of various aspects of reheating epoch.
Acknowledgements
We thank Kiwoon Choi, Suro Kim, Ryo Saito, Tomo Takahashi, Yuki Watanabe and Chul-Moon Yoo for helpful comments. This work is supported in part by Basic Science Research Program through the National Research Foundation of Korea (RS-2024-00336507). JG also acknowledges the Ewha Womans University Research Grant of 2025 (1-2025-0739-001-1). JG is grateful to the Asia Pacific Center for Theoretical Physics for hospitality during this work was under progress.